solve by the substitution method
3m+n=13
m-2n=23
what is the solution of the system

Answer by actuary(112)

Rearrange the first equation so that n=13-3m
SUBSTITUTE the rearranged equation into the second equation for n.
The second equation then becomes m-2(13-3m)=23
Use the DISTRIBUTIVE property so that this equation becomes
m-2*13+2*3m = 23 (REMEMBER - negative times a negative is a negative)
We now have m-26+6m = 23
The next step is to combine like terms
We now have 7m - 26 = 23
Add 26 to each side of the equation so that
7m = 49. Divide both sides of the equation by 7 so that m = 7
Substitute this value for m into the first equation to solve for n.
We now have 3*7+n=13
Therefore n = 13-3*7= 13-21 = -8
So m = 7 and n = -8